OpenStudy (anonymous):

Given that: 3 NO2(g) + H2O(l)  2 HNO3(aq) + NO(g), what quantity (L) of NO2 is required to completely react 125 grams of water. Assume that the gases are at standard temperature and pressure and hence 1.00 mole occupies 22.4 L of volume.

OpenStudy (mertsj):

\[125 g H _{2}O \times\frac{1 mole H _{2}O}{18 g H _{2}O}\times \frac{3 mole N _{2}O}{1mole H _{2}O}\times \frac{22.4 L N _{2}O}{1 mole N _{2}O}\]

OpenStudy (anonymous):

Thank you very much!!!

OpenStudy (mertsj):

yw